Use variation of parameters to solve the given nonhomogeneous system 4e-t X= x(t)- cie-t(-3,2) + c2e_4t(-1,1)
Answer #1

Complemervtay soln in veal eigen vauluo s Eigen vectors = (3-A)(-1-A) + 4 = atl) (A-3)+4 乀ー2-2バ.nal rs -20 -20 o 匕 only one s